OpenStudy (anonymous):

8 QUESTIONS PLZ HELP MEDAL AND FOLLOW

OpenStudy (anonymous):

whats the questions

OpenStudy (anonymous):

Question 1 (Essay Worth 4 points) Brenda is training to be a volunteer at her library. She wonders what the probability is that the first item she'll get to check out for a customer will be a video. So, on her first day of training, she tracks what items are checked out. The results are shown here in a tally chart. Brenda wrote tally marks for what type of item various customers check out. For books, there are seven groups of 5 tally marks and two individual marks. For videos, there are four groups of 5 tally marks and four individual marks. For magazines, there are three groups of 5 tally marks and four individual marks. Use Brenda's experimental data to estimate the probability that her first item to check out will be a video. Show and explain your work. Question 2 (Essay Worth 4 points) Shelly delivers the weekly local paper to neighborhoods in her town. House numbers are even on one side of the street and odd on the other. Shelly delivers an equal number of papers to both sides of the street. Although she always aims for the front doorstep, Shelly typically misses on three of the tosses on her route each week. Design and conduct a simulation to estimate the probability that next week, Shelly's three misses will all be at odd-numbered houses. Hint: You can set up the experiment using 3 coins to collect the data. Allow one side of the coin to represent Heads (evens) and one other side to represent Tails (odds). a) Explain clearly your design of the simulation, including choice of probability tool and description of a single trial. b) Conduct the simulation with twenty trials and record the results. c) Calculate the experimental probability that all 3 of Shelly's missed papers will be at odd-numbered houses. Question 3 (Multiple Choice Worth 2 points) Jaycee follows baseball closely. This season, his favorite professional player has a batting average of zero point two five zero, or twenty-five percent. This means that, on average, he gets one hit every four times he is at bat. Jaycee wants to know the likelihood that his favorite player will get at least two hits in the five times he'll be at bat. So Jaycee conducted a simulation to find out. He programmed a random number generator to give numbers between 1 and 4, with an outcome of 1 indicating a hit. His results are shown here. Trial 1: 3, 3, 1, 4, 2. Trial 2: 2, 2, 3, 3, 3. Trial 3: 3, 2, 2, 3, 3. Trial 4: 3, 4, 3, 3, 4. Trial 5: 1, 2, 4, 3, 4. Trial 6: 3, 3, 1, 4, 1. Trial 7: 4, 2, 1, 2 ,1. Trial 8: 1, 2, 4, 4, 2. Trial 9: 3, 2, 4, 2, 2. Trial ten: 3, 4, 1, 2, 4. Trial eleven: 3, 3, 1, 1, 1. Trial twelve: 4, 4, 2, 2, 2. Trial thirteen: 4, 2, 1, 2, 2. Trial fourteen: 4, 1, 1, 1, 3. Trial fifteen: 3, 2, 4, 2, 3. Trial sixteen: 1, 3, 3, 3, 3. Based on Jaycee's data, what is the probability that his favorite baseball player will get at least two hits in the next game, with five times at bat? twenty percent twenty-five percent seventy-five percent eighty percent Question 4 (Multiple Choice Worth 2 points) Kalil flipped a coin sixty times. He got thirty-four heads. How does this result compare to the number of heads you would have expected based on theoretical probability? It's less than expected. It's greater than expected. It's the same as expected. There is no expected value. Question 5 (Multiple Choice Worth 2 points) Which number of trials would best satisfy the need for a simulation to be something you could possibly do and provide meaningful data? 3 5 thirty-five three hundred fifty Question 6 (Multiple Choice Worth 2 points) A clothing store owner is experimenting with the layout of a clothing rack on the floor. One day, she kept track of the number of customers in the store who visited the rack of new clothing items. She found that sixty-two percent of her customers looked through that rack. The next day, she moved the rack of new items to a different location on the floor. Here's what she found: Total customers: seventy-five Number of customers looking at new items: sixty How did the movement of the rack affect the number of customers looking at the new items? It decreased by 2 percent It decreased by 8 percent It increased by thirteen percent It increased by eighteen percent Question 7 (Multiple Choice Worth 2 points) Which probability tool would not be an appropriate choice for a simulation that models an action that has four possible outcomes? a standard die a standard deck of cards a random number generator a spinner with equal sections labeled A, B, C, D Question 8 (Multiple Choice Worth 2 points) Which is the name given to a favorable outcome in an experiment? attempt probability tool success trial

OpenStudy (anonymous):

@MelissaHolmes @Whitemonsterbunny17

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@Confusionist

OpenStudy (anonymous):

@DSS

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@texaschic101 @Whitemonsterbunny17

OpenStudy (anonymous):

@OrangeMaster

OpenStudy (anonymous):

The reason that none of the people you've tagged has helped is because this is too much. Ask one question at a time, and the people will help if they want to. But, the point is that asking one at a time will get you much better results.

OpenStudy (anonymous):

1) 24/(37+24+19) ; 24/(37+24+19) = 1/3 2) 1/2 x 1/2 x 1/2 = 1/8 ; Its 1 eighth probability ; So 0.125; Or 12.5%;Or 1/8 3) divide the two and multiply by 100, so 4/16 = 0.25; then multiply by 100 to get 0.25*100 = 25%

OpenStudy (anonymous):

4) There is no expected value. because there is no way to tell if you are going to get heads or tails but you can guess which that would be 50/50 either right or wrong. 5) thirty-five 6) It increased by thirteen percent 7) a standard die 8) success hope this helps! :D